# Chen, Simchi-Levi, and Wang

### v2

@@ 8.1 Synthetic Data: Customer Lifetime Value
We revisit the example of customer lifetime value (Example 1.1), with a synthetic transition matrix detailed in Appendix 19.1. We assume that under the control policy, when the customer makes a purchase, the company will receive a net income of $30$ while the income in all other states is 00. Specifically, the transition matrix and the reward under the control policy are set as follows.
@@
    Local Treatment. The company plans to offer coupons whenever the customer’s recency is high (recency$>$2). We still consider five different candidate values. They uniformly decrease the rewards under states $s^{3},s^{4},s^{5}$ by $2,2.5,3,3.5,4$, respectively and increase the probability transiting to state $s^{1}$ by $0.10,0.11,0.12,0.13,0.14$, respectively.
We then run experiments for each candidate value under both cases treatment possible values, aiming to identify the optimal coupon value which are 4 and 2 . In each experiment, we run all three estimators over 1,000 independent trajectories and record their empirical mean and standard error. The experiment horizon is adjusted so that the IS estimator’s standard error is smaller than its empirical mean. In Figure 1, we present the simulation results.
@@
Variance reduction by IS and PI: In both examples, we observe a significant reduction in variance—by at least half—through the application of IS and PI. This illustrates the effectiveness of IS and PI in overcoming the lower bound of A/B testing by utilizing the local structure. Moreover, PI consistently outperforms IS in both examples, confirming its capability in further variance reduction.
@@ 8.2 A Real-World Sepsis Simulator
We run the simulations for $N=250$ patients over $100$ independent runs and then demonstrate the inference results of each estimators. We additionally compare two benchmarks, the Naive and Difference-in-Q (DQ) estimators, from Farias et al. (2022), which is originally designed for the general treatment policies on MDPs with two test arms. For the Naive estimators, we naively take the average on the rewards collected under the treatment policy and the control policy, ignoring the interference of state transitions. The DQ estimator strikes a balance between bias and variance under general and global treatment policies in MDPs. In Appendix 19.2, we generalize the DQ theory and provide the expression of DQ estimator for multiple test arms.
@@
Figure 2: Sepsis Simulator. Diamonds and caps represent the empirical mean and standard error respectively for each estimator. Left: Estimated ATE the best-performing arm among $n=8$ tested arms at patients $N=250$. Right: MSE of each estimator vs. $N$ the number of patients. IS, PI dominate over all experiment horizon.
@@
Variance Reduction of IS and PI: We observe a significant reduction in variance through the application of IS and PI while still obtaining negligible bias comparing with the A/B testing estimator. These methods also enjoy fast decay of MSE compared to all other estimators. This confirms the effectiveness of IS and PI and demonstrate the power of leveraging treatment structure to obtain striking efficiency improvement. Moreover, aligning with our intuition, PI consistently outperforms IS, confirming its capability in further variance reduction with more information.
Unbiasedness of IS and PI: Although the Naive estimator has smaller variance than IS and comparable variance to PI, its bias is non-negligible. Furthermore, the Naive estimator struggles to exclude the null effect, potentially resulting in insufficient statistical power. In contrast, the means of the IS and PI estimators are centered around the true ATE.
DQ exhibits bias-variance trade-offs: DQ demonstrates a notable bias-variance trade-off when compared to A/B testing and Naive estimators. DQ achieves variance reduction at the cost of introducing small bias, resulting in a faster decay of MSE. However, DQ fails to leverage the inherent local structures of the treatments, leading to significantly higher MSE compared to IS and PI. In Section 9.2, we introduce a novel estimator that combines DQ and IS concepts, demonstrating improved performance over the original DQ formulation.
@@ 9.1 Fairness And Welfare: The Monotonicity of SST
MDPs are complex systems with interconnected states. Therefore, out of fairness and welfare concern, it would be the best for the decision-makers to ensure that a new treatment improves long-term outcomes for customers across all states, rather than benefiting some at the expense of others. Fortunately, we have a theorem that guarantees the “monotonicity” of SST policies. This ensures that the treatment effect remains consistently positive (or negative) for all states before and after the intervention.
@@
The monotonicity of SST ensures that an improvement in any state will positively impact the long-term outcomes of all other states. Conversely, when the treatment effect is negative, all states except the crucial state will experience a welfare loss strictly less than $\gamma\Delta(s^{1})$, where $\gamma\in[0,1)$ is the discount factor. Additionally, the crucial state will always have the strongest treatment effect, aligning with our intuition that the crucial state is the primary target of the new treatment. Notably, this monotonicity theorem holds true regardless of the underlying Markov chain’s structure.
We provide two different proofs for Theorem 9.1 in Appendix 17. The first proof is based on an important observation that the treatment effect ${\bm{\Delta}}$ is an eigenvector of a special matrix associated with the transition probability under the control policy. By applying the Perron-Frobenius Theorem Meyer (2001) and carefully verifying the conditions, we arrive at the desired result. The second proof offers a novel perspective to characterize ${\bm{\Delta}}$ using the advantage function, derived from the well-knwon Performance Difference Lemma (Kakade and Langford 2002, Agarwal et al. 2022) in RL theory. We believe these techniques may be of independent interest to the RL literature.
9.2 Bias-Variance Trade-off
With general treatment, the seminal work by Farias et al. (2022) proposes a DQ estimator that trades a small bias for a striking variance reduction in average reward setting. In a follow-up work (Farias et al. 2023), they extend the idea of DQ to other reward settings including discounted reward used in this paper. In particular, we adapt their DQ theory to obtain the following estimator
$\widehat{\bm{\Delta}}^{\textrm{DQ}}\coloneqq\left({{\bm{I}}-\gamma\widehat{\bm{P}}_{1/2}}\right)^{-1}\left({\widehat{\bm{Q}}^{\pi_{1/2}}({\mathfrak{t}})-\widehat{\bm{Q}}^{\pi_{1/2}}({\mathfrak{c}})}\right)\,,$
where $\widehat{\bm{Q}}^{\pi_{1/2}}(a)=[\widehat{Q}^{\pi_{1/2}}(s^{i},a)]^{\top}_{i\in[K]}$ for $a\in\{{\mathfrak{t}},{\mathfrak{c}}\}$ fitted via TD Learning. We further observed that the concept of information sharing can be applied to this Q-function learning to reduce variance. This insight led us to develop a new estimator, which we call DQ-IS, taking advantage of two orthogonal ideas of DQ and IS.
In Appendix 19.3, we empirically implement both estimators for the SST setting. A surprising finding is that the DQ estimator shows minimal bias-variance trade-off for SST. Our empirical results indicate that both DQ and DQ-IS exhibit negligible bias, while their variance is close to that of A/B testing and IS respectively. We believe this phenomenon merits further investigation, as it has implications for both DQ theory and the broader literature on bias-variance trade-offs in RL.
9.3 Orthogonal Applications of Information Sharing
It is important to note that our information sharing approach is actually providing a more efficient way to utilize the observed trajectories by leveraging the local treatment structure. This concept of information sharing is not limited to discounted reward MDPs, but is also compatible with various other RL settings.
Average/Total Reward. Under different definitions of the value function, the information sharing idea is still applicable whenever the new treatment is localized. To be more specific, for the average reward MDPs, we can construct a similar plug-in model-based estimator following the corresponding Bellman Equation, ${\bm{V}}=({\bm{I}}-{\bm{P}})^{\sharp}{\bm{r}}$, where the notation $({\bm{I}}-{\bm{P}})^{\sharp}$ denotes the group inverse of the matrix $({\bm{I}}-{\bm{P}})$. Similarly, for the total reward, we have ${\bm{V}}=({\bm{I}}-{\bm{P}})^{-1}{\bm{r}}$ with ${\bm{P}}$ stands for the transition matrix of non-absorbing states. The estimations of ${\bm{P}}$ and ${\bm{r}}$ can still adopt the information sharing idea to reduce the variance.
@@ 9.3 Orthogonal Applications of Information Sharing
Episodic Setting. Another commonly used framework is the episodic MDP. In this scenario, multiple independent samples of the MDP can be obtained. For instance, the company in Example 1.1 might assume that certain customer groups share the same transition matrix. This allows for the observation of multiple trajectories, enabling the inference of the value function by aggregating data from these various episodes. Given this context, it’s natural to consider extending the concept of Information Sharing as a more sophisticated approach to data pooling in these scenarios. This extension could potentially lead to more efficient and accurate value function estimation by leveraging similarities across different episodes while accounting for their unique characteristics.
10 Conclusion
This paper systematically explores how treatment structure can be utilized to improve experimentation efficiency in MDPs. We first demonstrate that our A/B testing algorithm is efficient for general policies. Subsequently, we show how information can be shared among tests to reduce variance for local treatment policies, thereby overcoming the established lower bound.

### v3

@@ 7.1 Synthetic Data: Customer Lifetime Value
We revisit the example of customer lifetime value (Example 1.1), with a synthetic transition matrix detailed in Appendix 18.1. We assume that under the control policy, when the customer makes a purchase, the company will receive a net income of $30$ while the income in all other states is 0. Specifically, the transition matrix and the reward under the control policy are set as follows.
@@
    Local Treatment. The company plans to offer coupons whenever the customer’s recency is high (recency>>2). We still consider five different candidate values. They uniformly decrease the rewards under states $s^{3},s^{4},s^{5}$ by $2,2.5,3,3.5,4$, respectively and increase the probability transiting to state $s^{1}$ by $0.10,0.11,0.12,0.13,0.14$, respectively.
We additionally introduce another benchmark for comparison, called perfect information (PI), where the decision-makers have the access to the perfect information of the control arm. The PI estimators can always serve as the lower bound of the variance (i.e., the necessary variance). The detailed description of the PI estimators and their properties are provided in Appendix 15.
We run experiments for each candidate value under both cases treatment possible values, aiming to identify the optimal coupon value which are 4 and 2. In each experiment, we run all three estimators over 1,000 independent trajectories and record their empirical mean and standard error. The experiment horizon is adjusted so that the IS estimator’s standard error is smaller than its empirical mean. In Figure 1, we present the simulation results.
@@
Variance reduction by IS: In both examples, we observe a significant reduction in variance by at least half through the application of IS. The variance of IS is closer to the variance provided by PI. This illustrates the effectiveness of IS in overcoming the lower bound of original A/B testing by utilizing the local structure. Also, the more states the information can be shared, the closer the variance of IS estimator is to the PI estimator.
@@ 7.2 A Real-World Sepsis Simulator
We run the simulations for $N=250$ patients over $100$ independent runs and then demonstrate the inference results of each estimators. We additionally introduce two more benchmarks, the Naive and Difference-in-Q (DQ) estimators, from Farias et al. (2022), which is originally designed for the general treatment policies on MDPs with two test arms. For the Naive estimators, we naively take the average on the rewards collected under the treatment policy and the control policy, ignoring the interference of state transitions. The DQ estimator strikes a balance between bias and variance under general and global treatment policies in MDPs. In Appendix 18.2, we generalize the DQ theory and provide the expression of DQ estimator for multiple test arms.
@@
Figure 2: Sepsis Simulator. Diamonds and caps represent the empirical mean and standard error respectively for each estimator. Left: Estimated ATE the best-performing arm among $n=8$ tested arms at patients $N=250$. Right: MSE of each estimator vs. NN the number of patients. IS, PI dominate over all experiment horizon.
@@
Variance Reduction of IS: We observe a significant reduction in variance through the application of IS while still obtaining negligible bias comparing with the A/B testing estimator. The variance of IS is even close to the PI estimator. The IS and PI estimators also enjoy fast decay of MSE compared to all other estimators. This confirms the effectiveness of IS and demonstrates the power of leveraging treatment structure to obtain striking efficiency improvement.
Unbiasedness of IS: Although the Naive estimator has smaller variance than IS, its bias is non-negligible. Furthermore, the Naive estimator struggles to exclude the null effect, potentially resulting in insufficient statistical power. In contrast, the IS and PI estimators are centered around the true ATE.
DQ exhibits bias-variance trade-offs: DQ demonstrates a notable bias-variance trade-off when compared to A/B testing and Naive estimators. DQ achieves variance reduction at the cost of introducing small bias, resulting in a faster decay of MSE. However, DQ fails to leverage the inherent local structures of the treatments, leading to significantly higher MSE compared to IS and PI. As we have discussed that DQ belongs to the general differentiable estimators, we can expect that Is can improve performance over the original DQ formulation.
@@ 8.1 Fairness and Welfare: The Monotonicity of SST
MDPs are complex systems with interconnected states. Therefore, out of fairness and welfare concern, it would be the best for the decision-makers to ensure that a new treatment improves long-term outcomes for customers across all states, rather than benefiting some at the expense of others. Fortunately, we have a theorem that guarantees the “monotonicity” of SST policies.
@@
This ensures that the treatment effect remains consistently positive (or negative) for all states before and after the intervention. Furthermore, when the treatment effect is negative, all states except the crucial state will experience a welfare loss strictly less than $\gamma\Delta(s^{1})$, where $\gamma\in[0,1)$ is the discount factor. Notably, this monotonicity theorem holds true regardless of the underlying Markov chain’s structure.
Technically, we provide two different proofs for Theorem 8.1 in Appendix 16.2. The first proof is based on an important observation that the treatment effect $\bm{\Delta}$} is an eigenvector of a special matrix associated with the transition probability under the control policy. The second proof offers another perspective to characterize $\bm{\Delta}$} using the advantage function, derived from the well-knwon Performance Difference Lemma (Kakade and Langford 2002, Agarwal et al. 2022) in RL theory. We believe these techniques may be of independent interest to the RL literature.
8.2 Other Applications of Information Sharing
While this paper has mainly focused on A/B testing on a non-episodic MDP under the tabular setting, we believe it’s interesting directions to extend our framework to various other RL settings.
@@ 8.2 Other Applications of Information Sharing
Episodic Setting. Another commonly used framework is the episodic MDP. In this scenario, multiple independent samples of the MDP can be obtained. In section 7.2, each patients can be viewed as one such independent MDP. We thus empirically demonstrate that IS is also effective for episodic MDPs and brings about substantial variance reduction. We leave it to future work to derive rigorous theoretical guarantees for information sharing under episodic MDPs.
Policy Optimization. Most RL literature on finding optimal policy (e.g. Azar et al. 2017, Jin et al. 2018, Li et al. 2024) does not impose structural assumption on the action set. Given our modeling of treatment locality, a natural question is whether this structure can also be leveraged to improve the sample efficiency for policy optimization.
9 Concluding Remarks
This paper systematically explores how treatment structure can be utilized to improve the estimation of life-time effects. By exploiting treatment locality, our proposed estimators achieve substantial variance reduction, linear in the number of test arms, without sacrificing unbiasedness. We further show that a broad class of differentiable estimators can universally benefit from information sharing. These results not only establish new theoretical benchmarks but also highlight the practical value of local interventions for efficient experimentation in dynamic systems.